# Context as Memory: Scene-Consistent Interactive Long Video Generation with Memory Retrieval

### v1

@@ 8       Append $\mathbf{x}^{t}$ to $\mathcal{X}$ and $\mathbf{cam}^{t}$ to $\mathcal{C}$;
The second question is how to determine co-visibility between frames given their camera poses. We determine this by checking if there is an overlapping region between the fan-shaped areas corresponding to the Fields of View (FOV) of the two cameras. Specifically, since we restrict camera movement to the XY plane, we only need to consider the left and right rays shooting from each camera’s origin. By checking the intersection of these four rays from two cameras, we can quickly determine the FOV overlap as shown in Fig. 4. Additionally, we calculate the distance between the predicted frame’s camera and the calculated intersection points to eliminate cases where the cameras are too far apart (which typically indicates no actual overlap or very small overlap). This FOV overlap detection is not perfect, as it may fail in cases with occlusions. However, this method effectively reduces the number of candidate context frames.
Implementation details in training and inference. Assume the maximum number of retrieved context frames is $k$. During training, we read a long ground truth video (containing thousands of frames) and randomly select a segment as the sequence to be predicted. We then apply our Memory Retrieval method to select $k-1$ context frames from the remaining frames. The overlapping relationships between frames have been pre-computed, eliminating the need for repeated calculations. The first frame of the prediction sequence is also included as an additional context frame to ensure video continuity. Additionally, there is a $10\%$ probability during training that only the recent context frame is used, simulating the beginning of long video generation where no context frames are available. During inference, for each video segment to be predicted, we search $k-1$ context frames from the previously generated frames using FOV-based Memory Retrieval and add the most recently generated frame to the context. The training and inference procedures are outlined in Algorithm 1 and 2, respectively.
3.4 Data Collection
To validate our method, we require long video datasets with camera pose annotations. However, currently available datasets with camera pose information typically consist of short video clips [56, 2]. To obtain long-duration data with precise camera annotations, we utilized a simulation environment, specifically Unreal Engine 5. We generated randomized camera trajectories navigating through different scenes and rendered corresponding long videos. Our dataset comprises 100 videos of 7,601 frames each, featuring 12 distinct scene styles, with captions annotated by a multimodal LLM [48] every 77 frames. To simplify the problem while still effectively validating our method, we constrained the camera trajectory’s position changes to a 2D plane and limited rotation to only around the z-axis, which still presents sufficient complexity for camera trajectory control. Additional details about the dataset are provided in the supplementary materials.
4 Experiments
@@
DFoT [34] 17.63 0.4528 112.96 897.87 15.70 0.5102 121.18 919.75
FramePack [54] 17.20 0.4757 121.87 901.58 15.65 0.4947 131.59 974.52
@@
Table 1: Quantitative Comparison results. Due to learning abundant context, Context-as-Memory demonstrates the best memory and highest quality of generated videos. In contrast, DFoT [34] and FramePack [54], which can only utilize the most recent contexts, show relatively inferior performance, even worse than random context selection. This is because although random selection cannot guarantee to select useful information, on average it tends to obtain more information compared to methods that only learn from the most recent context.
Figure 5: Qualitative Comparison Results. Context-as-Memory demonstrates the best memory and the highest visual quality, indicating the effectiveness of sufficient context conditioning. Other methods exhibit scene inconsistency issues due to limited context utilization.
4.1 Experiment Settings
@@
Our method is implemented on an internal 1B-parameter pre-trained text-to-video Diffusion Transformer, developed for research purposes. The resolution of generated videos is $640\times 352$. The model supports generation of $77$-frame videos, with a temporal compression ratio of $4$ in the causal 3D VAE, resulting in $20$-frame video latents generation. We set the context size to $20$, meaning $20$ RGB frames are selected as context. Since these frames lack temporal continuity, they are individually compressed using the causal 3D VAE, also resulting in $20$ frames of video latents. The model was trained on our collected dataset for over $10,000$ iterations with a batch size of $64$ on $8$ NVIDIA A100 GPUs. During sampling, we employ Classifier-Free Guidance [15] for text prompts, with $50$ sampling steps.
@@
4.2 Comparison Results
In this section, we evaluate memory capabilities across baseline methods, SOTA approaches, and our Context-as-Memory. The compared methods include: (1) Single-frame context using the first frame; (2) Multi-frame context using the first frame plus random historical frames; (3) Diffusion Forcing Transformer (DFoT) [34], using a fixed-size window of most recent frames; (4) FramePack [54], which hierarchically compresses previous context into two frames, with each frame’s height or width halved compared to its predecessor. While theoretically supporting all historical frames, compression becomes impractical after several frames as latent size reduces to $1\times 1$. For fair comparison, all methods were implemented on our base model and dataset with identical training configurations and iterations. Results are presented in Tab. 1 and Fig. 5.
@@ 4.2 Comparison Results
Additionally, History Context Comparison proves more challenging than Ground Truth Comparison. Even with simple “rotate forward and rotate backward" trajectories, the performance gaps between methods are significant. DFoT and FramePack can only utilize the most recent context, causing them to continuously generate new content. Only by having access to global context and extracting useful relevant information from it can memory-aware new video generation be achieved.
4.3 Ablation Study
@@
Table 2: Ablation of Context Size. Larger context contain more useful information and lead to better memory, but also incur higher computational overhead, necessitating an optimal trade-off choice.
@@
FOV+Non-adj+Far-st 20.22 0.3003 18.11 0.3414
Table 3: Ablation of Memory Retrieval Strategy. The filtering methods of “FOV" and “Non-adj" (where only one frame from continuous frame sequences is selected as a candidate) effectively filter out useless and redundant information, leading to significant improvements in memory capability.
@@
We ablated different memory retrieval strategies to analyze their effects. “Random" refers to randomly selecting context; “FOV+Random" means first filtering using the FOV-based method, then randomly selecting from the remaining candidates; “Non-adj" means only one frame from continuous frame sequences will be selected as a candidate; “Far-st" means frames that are more distant in time or space are more likely to be selected. The results in Tab. 3 demonstrate the effectiveness of “FOV" and “Non-adj" methods in removing useless and redundant information, which significantly increases the probability of selecting useful context and thereby enhances memory capability. The impact of “Far-st" is relatively minor.
4.4 Open-Domain Results
Figure 6: Open-Domain Results. We collected open-domain images from the internet and used them as the first frame to generate subsequent long videos. Under the trajectory of “rotate away and rotate back," even when generating new content, it still demonstrates good memory capability.
Due to our diverse training dataset and the various visual priors learned by our base model during pre-training, our method has the potential to generalize to open-domain scenarios not present in the training set. We selected images of different styles from the internet and used them as the first frame to generate long videos. We validated using the trajectory of “rotate away and rotate back," which is suitable for verifying memory consistency in generated content. Results in Fig. 6 demonstrate that our method indeed possesses good memory capability in open-domain scenarios.
5 Conclusion

### v2

@@ 8   Append $\mathbf{x}^{t}$ to $\mathcal{X}$ and $\mathbf{cam}^{t}$ to $\mathcal{C}$;
3.4. Data Collection
To validate our method, we require long video datasets with camera pose annotations. However, currently available datasets with camera pose information typically consist of short video clips (Zhou et al., 2018; Bai et al., 2025). To obtain long-duration data with precise camera annotations, we utilized a simulation environment, specifically Unreal Engine 5. We generated randomized camera trajectories navigating through different scenes and rendered corresponding long videos. Our dataset comprises 100 videos of 7,601 frames each, featuring 12 distinct scene styles, with captions annotated by a multimodal LLM (Yao et al., 2024) every 77 frames. To simplify the problem while still effectively validating our method, we constrained the camera trajectory’s position changes to a 2D plane and limited rotation to only around the z-axis, which still presents sufficient complexity for camera trajectory control. Additional details about the dataset are provided in the supplementary materials.
4. Experiments
Figure 5. Qualitative Comparison Results. Among them, Context-as-Memory demonstrated the best memory capabilities and the highest visual quality, indicating the effectiveness of sufficient context information conditioning. Other methods exhibit scene inconsistency issues due to limited context utilization.
@@ 4. Experiments
DFoT (Song et al., 2025) 17.63 0.4528 112.96 897.87 15.70 0.5102 121.18 919.75
FramePack (Zhang and Agrawala, 2025) 17.20 0.4757 121.87 901.58 15.65 0.4947 131.59 974.52
@@
Table 1. Quantitative Comparison results. Due to learning abundant context, Context-as-Memory demonstrates the best memory capabilities and highest quality of generated videos. In contrast, DFoT (Song et al., 2025) and FramePack (Zhang and Agrawala, 2025), which can only utilize the most recent contexts, show relatively inferior performance, even worse than random context selection. This is because although random selection cannot guarantee the selection of useful information, on average it tends to obtain more information compared to methods that only learn from the most recent context.
4.1. Experiment Settings
@@
Our method is implemented on an internal 1B-parameter pre-trained text-to-video Diffusion Transformer, developed for research purposes. The resolution of generated videos is $640\times 352$. The model supports generation of $77$-frame videos, with a temporal compression ratio of $4$ in the causal 3D VAE, resulting in $20$-frame video latents generation. We set the context size to $20$, meaning $20$ RGB frames are selected as context. Since these frames lack temporal continuity, they are individually compressed using the causal 3D VAE, also resulting in $20$ frames of video latents. The model was trained on our collected dataset for over $10,000$ iterations with a batch size of $64$ on $8$ NVIDIA A100 GPUs. During sampling, we employ Classifier-Free Guidance (Ho and Salimans, 2022) for text prompts, with $50$ sampling steps.
@@
4.2. Comparison Results
In this section, we evaluate video generation memory capabilities across baseline methods, SOTA approaches, and our Context-as-Memory. The compared methods include: (1) Single-frame context using the first frame; (2) Multi-frame context using the first frame plus random historical frames; (3) Diffusion Forcing Transformer (DFoT) (Song et al., 2025), using a fixed-size window of most recent frames; (4) FramePack (Zhang and Agrawala, 2025), which hierarchically compresses previous context into two frames, with each frame’s height or width halved compared to its predecessor. While theoretically supporting all historical frames, compression becomes impractical after several frames as latent size reduces to $1\times 1$. For fair comparison, all methods were implemented on our base model and dataset with identical training configurations and iterations. Results are presented in Table 1 and Figure 5.
@@ 4.2. Comparison Results
Additionally, History Context Comparison proves more challenging than Ground Truth Comparison. Even with simple ”rotate forward and rotate backward” trajectories, the performance gaps between methods are significant. DFoT and FramePack can only utilize the most recent context, causing them to continuously generate new content. Only by having access to global context and extracting useful relevant information from it can memory-aware new video generation be achieved.
4.3. Ablation Study
@@
Table 2. Ablation of Context Size. Larger context sizes contain more useful information and lead to better memory capability, but also incur higher computational overhead, necessitating an optimal trade-off choice.
@@
FOV+Non-adj+Far-space-time 20.22 0.3003 18.11 0.3414
Table 3. Ablation of Memory Retrieval Strategy. The filtering methods of ”FOV” and ”Non-adj” (where only one frame from continuous frame sequences is selected as a candidate) effectively filter out useless and redundant information, leading to significant improvements in memory capability.
@@
We ablated different memory retrieval strategies to analyze their effects. ”Random” refers to randomly selecting context; ”FOV+Random” means first filtering using the FOV-based method, then randomly selecting from the remaining candidates; ”Non-adj” means only one frame from continuous frame sequences will be selected as a candidate; ”far-space-time” means frames that are more distant in time or space are more likely to be selected. The results in Tab. 3 demonstrate the effectiveness of ”FOV” and ”Non-adj” methods in removing useless and redundant information, which significantly increases the probability of selecting useful context and thereby enhances memory capability. The impact of ”Far-space-time” is relatively minor.
4.4. Open-Domain Results
Figure 6. Open-Domain Results. We collected open-domain images from the internet and used them as the first frame to generate subsequent long videos. Under the trajectory of ”rotate away and rotate back,” even when generating new content, it still demonstrates good memory capability.
Due to our diverse training dataset and the various visual priors learned by our base model during pre-training, our method has the potential to generalize to open-domain scenarios not present in the training set. We selected images of different styles from the internet and used them as the first frame to generate long videos. We validated using the trajectory of ”rotate away and rotate back,” which is suitable for verifying memory consistency in generated content. Results in Fig. 6 demonstrate that our method indeed possesses good memory capability in open-domain scenarios.
5. Conclusion
@@
Appendix A Introduction of the Base Text-to-Video Generation Model
We use a transformer-based latent diffusion model as the base T2V generation model, as illustrated in Fig. 7. We employ a 3D-VAE to transform videos from the pixel space to a latent space, upon which we construct a transformer-based video diffusion model. Unlike previous models that rely on UNets or transformers, which typically incorporate an additional 1D temporal attention module for video generation, such spatially-temporally separated designs do not yield optimal results. We replace the 1D temporal attention with 3D self-attention, enabling the model to effectively perceive and process spatiotemporal tokens, thereby achieving a high-quality and coherent video generation model. Specifically, before each attention or feed-forward network (FFN) module, we map the timestep to a scale, thereby applying RMSNorm to the spatiotemporal tokens.
Appendix B Details of Collected Dataset
In this section, we provide a detailed description of the rendered dataset used to train our model.
3D Environments
We collect 12 different 3D environments assets from https://www.fab.com/. To minimize the domain gap between rendered data and real-world videos, we primarily select visually realistic 3D scenes, while choosing a few stylized or surreal 3D scenes as a supplement. To ensure data diversity, the selected scenes cover a variety of indoor and outdoor settings, such as city streets, shopping malls, and the countryside.
Camera Trajectories
To create data that roam within a scene, we employ smoothed polylines as camera trajectories. Specifically, we begin by randomly sampling coordinate points in the 3D scene to serve as the endpoints of the polyline, and then generate B-spline curves from these points. To ensure smooth camera movement without abrupt speed changes or rotations, we limit the camera’s movement distance to the range of [3m, 6m] for each 77-frame video segment and restrict the rotation angle within the xy-plane to less than 60 degrees.
Upon completing the 3D scene collection and trajectory design, we utilized Unreal Engine 5 to batch-render 100 long videos for training. Each video features 7,601 frames (30 fps) of continuous camera movement. Additionally, we record the camera’s extrinsic and intrinsic parameters for each frame. The camera is configured with a focal length of 24mm, an aperture of 10, and a field of view (FOV) of 52.67 degrees.
Appendix C Additional Open-Domain Results
In Fig. 8 and Fig. 9, we present additional open-domain results. Using diverse images collected from the internet as initial frames, we demonstrate long video generation with ”rotate away and rotate back” trajectories. These source images, representing various styles and scenes, can be found in the provided Data.
Our method achieves generalization capability in open-domain scenarios due to two main factors: (1) Training on diverse scenes enables the model to develop generalizable context utilization skills; (2) The pre-trained base model possesses strong generative priors from exposure to various data types during pre-training.
However, our method still faces significant limitations in open-domain generalization that require future research: (1) The 1B-parameter base model’s capabilities are insufficient, only showing good results on simple trajectories. For complex trajectories, the base model struggles to generate high-quality content from the initial frame, leading to unacceptable error accumulation in long video generation. Validating our approach with larger-scale base models remains a future research direction. (2) The method cannot yet support more complex, diverse, and dynamic long-term scene exploration in open-domain settings. Our ideal goal is to enable free, extended navigation from any given image while maintaining memory consistency. This is a challenging objective, though the ”context as memory” concept shows promise.
Figure 8. Open-Domain Results.
Figure 9. Open-Domain Results.